OpenStudy (lilkg77):

which of these points lies on the line described by the equation below? y-8=3(x+6) A. (8, -6) B. (6, -8) c. (-8, 6) D. (-6, 8)

OpenStudy (anonymous):

what do u think?

OpenStudy (lilkg77):

Im not sure

Directrix (directrix):

Test the options to see if the coordinates given make the equation true. To help you get started, I will check the first option. Is A. (8, -6) a point on the line with equation: y-8=3(x+6) x = 8 and y = -6 y-8=3(x+6) -6 - 8 = 3 * ( 8 + 6 ) ? -14 = 3 * 14 NO The point in option A is not on the line. Test option B @lilkg77

OpenStudy (lilkg77):

How do i do that without a graph?

OpenStudy (anonymous):

find the slope

OpenStudy (anonymous):

you know (-6,8) is already a point on the line from the equation

OpenStudy (anonymous):

the y intercept is (0,26)

OpenStudy (anonymous):

oh pfft the answer is d ! xD

OpenStudy (anonymous):

You should study point-slope form before doing this, it is actually quite simple

OpenStudy (lilkg77):

Lol thanks man @TheSavior

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

i also say it's D.